A pharmacist wants to prepare a 60liter mixture that is 40% boola using three different kinds of boola. The first concentration is 15% hula boola, the second is 35 % mula boola, and the third is 55 % gula boola. Because of the amounts of the boola solutions available, the pharmacist wants to use twice as much of the mula solution as the gula solution. If no chemical reactions among three different boolas take place, how much of each solution should be used?
Answer by ankor@dixie-net.com(22740)

A pharmacist wants to prepare a 60liter mixture that is 40% boola using three different kinds of boola.
3 substances x + y + z = 60
:
The first concentration is 15% hula boola, the second is 35 % mula boola, and the third is 55 % gula boola.
.15x; .35y; .55z
:
Because of the amounts of the boola solutions available, the pharmacist wants to use twice as much of the mula solution as the gula solution.
y = 2z
therefore
x + 2z + z = 60
x + 3z = 60
x = (60-3z)
:
If no chemical reactions among three different boolas take place, how much of each solution should
:
.15x + .35y + .55z = .40(60)
:
x = (60-3z) and y = 2z
.15(60-3z) + .35(2z) + .55z = 24
9 - .45z + .70z + .55z = 24
.80z = 24 - 9
.80z = 15
z = 15/.8
z = 18.75 liters of g.b
then
x = 60-3(18.57)
x = 60 - 56.25
x = 3.75 liters of h.b
and
y = 2(18.75)
y = 37.5 liters of m.b.
:
:
see if this mess checks out
.15(3.75) + .35(37.5) + .55(18.75) = .40(60)
.5625 + 13.125 + 10.3125 = 24, it does!